OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4.

OpenStudy (anonymous):

The first thing you want to do is put the equation in slope intercept form. Slope intercept form: y = mx+b so you would subtract 3x from the equation and put it in the other side. -3y=4-3x You don't want the y to have a number in front of it, so divide the entire side of the second equation by -3 to get rid of the -3. y = -4/3 + 1x or y = x - 4/3 The slope is 1 and the y intercept is -4/3

OpenStudy (anonymous):

wait let me check though

OpenStudy (anonymous):

Yeah, it's right

OpenStudy (anonymous):

oh alright thanks